Suppose that a rectangle has a perimeter of 22 meters.
Express the area A(x) of the rectangle in terms of the length
x of one of its sides.

Answer by RAY100(1637)

Perimeter = 2Length + 2 Width =22 (given)
.
divide by 2 thruout,,,,L + W = 11,,,or L=11-W
.
Area = Length * Width,,,,,,if we let w=x, and substitute
.
A=(11-W) * (W) = (11-x) *(x) = 11x -x^2
.
or,,,,A= 11x-x^2
.
check,,,let L=10,,W=1,,,Per = 2(10)+ 2(1) = 22,,,,ok
.
A=1*10=10,,,,,,A= 11(1)-(1)^2 = 10,,,,ok
.
.
For L=9...w=2,,,,per =22 ok,,,,,A=9*2=18,,,A= 11(2)-2^ = 18,,,ok
.